OpenStudy (anonymous):

Hi I am having trouble with this type of partial fraction decompisition. 2/(x^3 + x^2)

OpenStudy (anonymous):

I think I need A B and C...the needing 3 parts is what is confusing me.

zepdrix (zepdrix):

\[\Large \frac{2}{x^3+x^2}\quad=\quad \frac{2}{x^2(x+1)}\]Hmm so I guess that's as far as we want to factor it. having a bit of trouble getting the proper setup?

zepdrix (zepdrix):

We have a repeated linear factor `x` in the bottom, so we have to do something fancy for that one.\[\Large \cfrac{2}{x^2(x+1)} \quad=\quad \cfrac{A}{x}+\cfrac{B}{x^2}+\cfrac{C}{x+1}\]Mmmmm yah, I think this is the setup that we want.

OpenStudy (ranga):

@zepdrix's setup works nicely!

zepdrix (zepdrix):

Partial fractions have never made a ton of sense to me so I can't think of the proper way to explain this... Notice the bottom you can write as x*x*(x+1) So we have a repeated linear factor of x. There is a special way we deal with those types of things. I would recommend googling `partial fraction decomposition repeated factors` if you want to get some better understanding of that rule. Hmm, so do you understand how to solve the problem from that setup?

OpenStudy (anonymous):

so A/x + B/x^2 +C/x+1) ?

OpenStudy (anonymous):

ok thank you!